factorise by using factor theorum :- y^3-7y+6 (please show its steps also)
Answer by lwsshak3(11628)

factorise by using factor theorum :- y^3-7y+6 (please show its steps also)
**
f(y)=y^3-7y+6
Use rational roots theorem to solve: (with synthetic division)
...0..|...-1......0......-7.......6............
...1..|.....1......1......-6......0 (1 is a root)
f(y)=(y-1)(y^2+y-6)
f(y)=(y-1)(y+3)(y-2)